d) Wavenumber of the transition 211 132 e) Wavenumber of the transition 000001
8. Calculate the frequency of transitions using the following informations: Asymmetric, bending and symmetric modes vi, v2, and va for hydrogen eyanide, HCN molecule are 3311 cm1, 713 em' and 2097 cm respectively. a) Wavenumber of the transition 000-040 b) Wavenumber of the transition: 111222 c) wavenumber of the transition 011 → 022
d) Wavenumber of the transition 211 132 e) Wavenumber of the transition 000001
Answer #1

Answer:

\nu1 = 3311cm-1\nu2 = 713 cm-1\nu3 = 2097 cm-1

energy required for going from a,b,c to x,y,z

= ( E(x,y,z) - E(a,b,c) )

= h\nu1(x + 1/2) + h\nu2(y + 1/2) + h\nu3(z + 1/2) - ( h\nu1(a + 1/2) + h\nu2 (b + 1/2) + h\nu3(z + 1/2) )

= h (\nu1(x - a) + \nu2(y - b) + \nu3(z - c))

So the wavenumbers corrsponding to above energy = E/h = \nu1(x - a) + \nu2(y - b) + \nu3(z - c)

So the answer for the above questions is

a) 000 --> 040

E( in cm-1) = (\nu1(0 - 0) + \nu2(4 - 0) + \nu3(0 - 0)) = 4\nu2 = 4 x 713 cm-1 = 2852 cm-1

b) 111 --> 222

E( in cm-1) = (\nu1(2 - 1) + \nu2(2 - 1) + \nu3(2 - 1)) = \nu1 + \nu2 + \nu3 = (3311 + 713 + 2097) cm-1 = 6121 cm-1

c) 011 --> 022

E( in cm-1) = (\nu1(0 - 0) + \nu2(0 - 0) + \nu3(2 - 1)) = \nu2 + \nu3 = (713 + 2097) cm-1 = 2810 cm-1

d) 211 --> 132

E( in cm-1) = (\nu1(1 - 2) + \nu2(3 - 1) + \nu3(2 - 1)) = -\nu1 + 2\nu2 + \nu3 = (-3311 + (2 x 713) + 2097) cm-1 = 212 cm-1

d) 000 --> 001

E( in cm-1) = (\nu1(0 - 0) + \nu2(0 - 0) + \nu3(1 - 0)) = 0\nu1 + 0\nu2 + \nu3 = \nu3 = 2097 cm-1